OpenStudy (anonymous):

a drugist has one solution that is 10% iodine and another that is 50% iodine. How much of each should he use to get 100mL of a mixture that is 20% iodine?

jimthompson5910 (jim_thompson5910):

let x = amount of 10% iodine and y = amount of 50% iodine

jimthompson5910 (jim_thompson5910):

since you make a final mix of 100 mL, we know that x+y = 100 and you can solve for y to get y = 100 - x

jimthompson5910 (jim_thompson5910):

you want to make a final mix of 20% iodine, so you want 100*0.2 = 20 mL of pure iodine there are 0.1x mL of pure iodine from the first solution and 0.5y mL of pure iodine from the second

jimthompson5910 (jim_thompson5910):

so 0.1x + 0.5y = 20

OpenStudy (anonymous):

is tht the answer

jimthompson5910 (jim_thompson5910):

no, it helps you find the answer

jimthompson5910 (jim_thompson5910):

you need to solve for x and y

OpenStudy (anonymous):

can u go futher for me

jimthompson5910 (jim_thompson5910):

0.1x + 0.5y = 20 0.1x + 0.5(100 - x) = 20 now solve for x

OpenStudy (anonymous):

distribute?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

0.3?

jimthompson5910 (jim_thompson5910):

0.1x + 0.5y = 20 0.1x + 0.5(100 - x) = 20 0.1x + 50 - 0.5x = 20 -0.4x + 50 = 20 Did you get this far?

OpenStudy (anonymous):

0.1+0.5(100-x)=20 0.6x(100-x)=20 60x-x=20 +x +x ------------ 60x=20 --- --- 60 60 x=0.3

jimthompson5910 (jim_thompson5910):

0.1x + 0.5y = 20 0.1x + 0.5(100 - x) = 20 0.1x + 50 - 0.5x = 20 -0.4x + 50 = 20 -0.4x = 20-50 -0.4x = -30 x = -30/(-0.4) x = 75

OpenStudy (anonymous):

nice work jim

jimthompson5910 (jim_thompson5910):

thanks